OpenStudy (anonymous):

Quadrilateral ABCD with vertices A(4, 3), B(4, -2), C(-4, -2) and D(-4, 3) is a rectangle, find the length of the diagonals. A. in comments B. 5 C. 8 D. in comments

OpenStudy (anonymous):

A)\[\sqrt{89}\] B) \[\sqrt{39}\]

OpenStudy (anonymous):

Ok so this forms a shape that can be divided into two to form two right triangles.

OpenStudy (anonymous):

So since we have the coordinates, you can use the distance formula to find the distance between the points.

OpenStudy (anonymous):

Distance formula is \[\sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

Can you go on from there?

OpenStudy (anonymous):

yes thank you :)